OpenStudy (istim):

Determine the exact equation of the function given: f(x)=k(x-1)(x+2) and goes thru point (2,-8)

OpenStudy (istim):

I'm looking through my notes, and it gives info SO close to what I want...

Directrix (directrix):

-8 =k(2-1)(2+2) -8 = k (1)(4) -8= 4k k=-2

OpenStudy (istim):

So all I do was sub in values? Ok. Thank you.

Directrix (directrix):

Yes. You may need to multiply out the factors and express the answer in polynomial form. Check the directions.

OpenStudy (istim):

"Multiply out"? And what do you mean by polynomial form?

OpenStudy (chris):

He's saying that polynormial form is http://en.wikipedia.org/wiki/Polynomial - e.g. ax^2 + bx + c - basically the opposite of factored. really depends on how they want the final format of the equation

OpenStudy (chris):

e.g. if k = -2, then one way to write the equation is f(x) = -2(x-1)(x+2)

OpenStudy (chris):

another way is to FOIL (x-1)(x+2) to put the final equation in polynomial form

OpenStudy (istim):

Ok. Thank you Chris.

OpenStudy (chris):

Basically it's up to convention - they are equivalent

OpenStudy (chris):

np!